OpenStudy (anonymous):

sqrt x+3-2=2 solve for x

OpenStudy (anonymous):

\[\sqrt{x+3-2=2}\]

OpenStudy (anonymous):

help ? @KayleeNeedsHelp @mathmath333 @Maria195 @Tazmaniadevil @love10129151 @plzhelp!?!? @ilovewolf @WWWHHHAAATTT? @micahm @Kit_Kat_Nat_<3 @Catlover5925

OpenStudy (anonymous):

@freckles @camper4834

OpenStudy (anonymous):

@sangya21 @kah.x

OpenStudy (wwwhhhaaattt?):

@kdog55 What do you think it is?

OpenStudy (wwwhhhaaattt?):

Ok, Try

OpenStudy (wwwhhhaaattt?):

OK, TRY AGAIN, RIGHT HERE

OpenStudy (wwwhhhaaattt?):

No, im trying to help

OpenStudy (anonymous):

anybody else can help , ill give medals @Pi123 @wade123 @wwhitlock @undeadknight26 @jigglypuff314 @iambatman

OpenStudy (wwwhhhaaattt?):

Like i said, Try

OpenStudy (wwwhhhaaattt?):

If you try i will show you how

OpenStudy (anonymous):

i tried that , could anybosy just give me and answer , this is my last hw question

OpenStudy (wwwhhhaaattt?):

Giving Answers is agianst CoC

OpenStudy (wwwhhhaaattt?):

Just Try

OpenStudy (anonymous):

bye @WWWHHHAAATTT?

OpenStudy (wwwhhhaaattt?):

Im Waiting for you to try

OpenStudy (wwhitlock):

\[\sqrt{x+3-2}=2\] \[\sqrt{x+1}=2\] x=3